# Pauli Hamiltonians with an Aharonov-Bohm flux

### v1

@@ 2.3. Spectral and scattering properties
Making reference to the quadratic form parametrization $H_{\mathrm{P}}^{(\beta)}$ of the Pauli Hamiltonian, using the bijection in Proposition 2.8 and noting that $\Lambda(0)=0$ (see (2.12)), we get
@@ 2.4. Symmetries and connection with the Dirac operator
$\|H_{\mathrm{D}}\bm{\psi}\|^{2}=\left\langle\bm{\psi}\left|H_{\mathrm{P}}\right|\bm{\psi}\right\rangle=Q^{(\mathrm{F})}_{P}(\bm{\psi})\,,$
@@
$\mathscr{D}\big{(}\overline{H}_{\mathrm{D}}\big{)}=\mathscr{D}\big{[}Q^{(\mathrm{F})}_{P}\big{]}\,.$ (2.59)
@@
For any $\alpha\in(0,1)$, the symmetric operator $H_{\mathrm{D}}$ admits a one-parameter family of self-adjoint extensions $H_{\mathrm{D}}^{\gamma}$, parametrized by $\gamma\in[0,2\pi)$ and given by
$\mathscr{D}(H_{\mathrm{D}}^{\gamma})=\left\{\bm{\psi}=\bm{\phi}+\mu\left(\bm{\xi}_{+}+e^{i\gamma}\bm{\xi}_{-}\right)\,\Big{|}\,\bm{\phi}\in\mathscr{D}[Q^{(\mathrm{F})}_{P}]\,,\;\mu\in\mathbb{C}\right\},$ (2.61)
$H_{\mathrm{D}}^{\gamma}\bm{\psi}=H_{\mathrm{D}}\bm{\phi}+i\mu(\bm{\xi}_{+}-i\bm{\xi}_{-})\,.$ (2.62)
Alternatively, as illustrated in [Pe06], the domain of the extensions $H_{\mathrm{D}}^{\gamma}$ can be described in terms of boundary conditions at the origin, as follows: given $\gamma\in[0,2\pi)$, one can define the linear functionals $c^{\uparrow,\downarrow}_{-\alpha}$, $c^{\uparrow,\downarrow}_{\alpha-1}$ on $\mathscr{D}(H_{\mathrm{D}}^{\gamma})$ as
$c^{s}_{-\alpha}(\bm{\psi})=\lim_{r\to 0^{+}}r^{\alpha}\,\langle\psi_{s}\rangle=\tfrac{1}{2\pi}\lim_{r\to 0^{+}}r^{\alpha}\int^{2\pi}_{0}\mathrm{d}\theta\,\psi_{s}(re^{i\theta})\,,$ (2.63)
$c^{s}_{\alpha-1}(\bm{\psi})=\lim_{r\to 0^{+}}r^{1-\alpha}\,\langle e^{i\theta}\psi_{s}\rangle=\tfrac{1}{2\pi}\lim_{r\to 0^{+}}r^{1-\alpha}\int^{2\pi}_{0}\mathrm{d}\theta\,e^{i\theta}\,\psi_{s}(re^{i\theta})\,,$ (2.64)
for $s\in\left\{\uparrow,\downarrow\right\}$. Observe that, for a given $\bm{\psi}\in\mathscr{D}(H_{\mathrm{D}}^{\gamma})$, we have a decomposition $\bm{\psi}=\bm{\phi}+\mu(\bm{\xi}_{+}+e^{i\gamma}\bm{\xi}_{-})$ as in Proposition 2.22, so that $\bm{\phi}$ vanishes at the origin, thus having zero boundary value. The asymptotics of Bessel functions $K_{\nu}$ at the origin [GR07] then easily give
@@
where $\Gamma(z)$ is the Euler gamma function. We can then equivalently describe the domain of $H_{\mathrm{D}}^{\gamma}$ as
$\mathscr{D}(H_{\mathrm{D}}^{\gamma})=\left\{\bm{\psi}\in L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})\,\Big{|}\,H_{\mathrm{D}}\bm{\psi}\in L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})\,,\;c^{\uparrow}_{\alpha-1}(\bm{\psi})=i\cot(\gamma/2)\,\tfrac{2^{1-2\alpha}\Gamma(1-\alpha)}{\Gamma(\alpha)}\,c^{\downarrow}_{-\alpha}(\bm{\psi})\,,\right.\\ \left.c^{\uparrow}_{-\alpha}(\bm{\psi})=0\,,c^{\downarrow}_{\alpha-1}(\bm{\psi})=0\right\}.$ (2.67)
We have already pointed out that the symmetric operators $H_{\mathrm{P}}$ and $H_{\mathrm{D}}$ defined on $C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{0\};\mathbb{C}^{2})$, are such that $H_{\mathrm{D}}^{2}=H_{\mathrm{P}}$. Now we investigate whether the same property holds for the the respective self-adjoint extensions. More precisely, we aim at determining which extension $H_{\mathrm{P}}^{(\beta)}$ is the square of an extension $H_{\mathrm{D}}^{\gamma}$, if any. It turns out that the answer to this question is actually trivial, since the square of all self-adjoint realizations of the Dirac operator coincides with the Fridriechs extension of $H_{\mathrm{P}}$. As a byproduct of the proof, we are also able to classify all the extensions $H_{\mathrm{P}}^{(\beta)}$ such that $\mathscr{D}(H_{\mathrm{P}}^{(\beta)})\subseteq\mathscr{D}(H_{\mathrm{D}}^{\gamma})$ for some $\gamma\in[0,2\pi)$, i.e., both the first and second order operators are simultaneously well-posed, although in general the latter is not the square of the former.
@@
$\left(H_{\mathrm{D}}^{\gamma}\right)^{2}=H_{\mathrm{P}}^{(\mathrm{F})}\,.$ (2.68)
Furthermore, $\mathscr{D}(H_{\mathrm{P}}^{(\beta)})\subseteq\mathscr{D}(H_{\mathrm{D}}^{\gamma})$, for some $\gamma\in[0,2\pi)$, if and only if, for all $\lambda>0$ large enough,
@@ 3.1. Quadratic forms
$g^{(\ell)}_{\lambda}\!(r,\theta)=\left[\tfrac{\Gamma(|\ell+\alpha|)}{2^{1-|\ell+\alpha|}}\,{1\over r^{|\ell+\alpha|}}+\tfrac{\Gamma(\!-|\ell+\alpha|)}{2^{1+|\ell+\alpha|}}\,\lambda^{2|\ell+\alpha|}\,r^{|\ell+\alpha|}+\mathcal{O}\big{(}r^{2-|\ell+\alpha|}\big{)}\right]\tfrac{e^{i\ell\theta}}{\sqrt{2\pi}}\,,\quad\mbox{for\, $r\to 0^{+}$}.$ (3.2)
@@ 3.2. Krein’s theory
$\big{|}I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z}\,(r\wedge r^{\prime})\big{)}\big{|}\leqslant\left(\!\tfrac{\sqrt{|z|}}{\Im\sqrt{z}}\right)^{\!|\ell+\alpha|}I_{|\ell+\alpha|}\big{(}\Im\sqrt{z}\,(r\wedge r^{\prime})\big{)}\,;$ (3.5)
$\big{|}K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z}\,(r\land r^{\prime})\big{)}\big{|}\leqslant K_{|\ell+\alpha|}\big{(}\Im\sqrt{z}\,(r\wedge r^{\prime})\big{)}\,.$ (3.6)
@@
$\tau_{s}^{(\ell)}R_{\mathrm{S}}^{(\mathrm{F})}(z^{*})\,\bm{\psi}\\ =\lim_{r\to 0^{+}}\!\tfrac{2^{|\ell+\alpha|-1}\,\Gamma\big{(}|\ell+\alpha|\big{)}}{r^{|\ell+\alpha|}}\;\big{(}|\ell+\alpha|+r\,\partial_{r}\big{)}\!\int^{\infty}_{0}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\,I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,(r\wedge r^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,(r\land r^{\prime})\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\\ =2^{|\ell+\alpha|-1}\,\Gamma\big{(}|\ell+\alpha|\big{)}\,i\sqrt{z^{*}}\lim_{r\to 0^{+}}\!\,r^{1-|\ell+\alpha|}\,\bigg{[}K_{|\ell+\alpha|-1}\big{(}\!-i\sqrt{z^{*}}\,r\big{)}\!\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\\ -I_{|\ell+\alpha|-1}\big{(}\!-i\sqrt{z^{*}}\,r\big{)}\!\int^{\infty}_{r}\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\,K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\bigg{]}\,.$
@@
$\left|\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\right|\leqslant\big{\|}\psi_{s}^{(\ell)}\big{\|}_{L^{2}(\mathbb{R}^{+},\,r\,dr)}\left(\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,\big{|}I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\big{|}^{2}\right)^{\!\!1/2}\\ \leqslant\left(\!\tfrac{\sqrt{|z|}}{\Im\sqrt{z^{*}}}\right)^{\!|\ell+\alpha|}\big{\|}\psi_{s}^{(\ell)}\big{\|}_{L^{2}(\mathbb{R}^{+},\,r\,\mathrm{d}r)}\left(\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,I^{2}_{|\ell+\alpha|}\big{(}\Im\sqrt{z^{*}}\,(r\wedge r^{\prime})\big{)}\!\right)^{\!\!1/2}\leqslant C\,\|\bm{\psi}\|_{2}\;r^{1+|\ell+\alpha|}\xrightarrow[r\to 0^{+}]{}0\,,$
@@
$\tau_{s}^{(\ell)}R_{\mathrm{S}}^{(\mathrm{F})}(z^{*})\,\bm{\psi}=\int_{0}^{\infty}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\!\int_{0}^{2\pi}\!\!\!\!\mathrm{d}\theta^{\prime}\,\big{(}\!-i\sqrt{z^{*}}\big{)}^{|\ell+\alpha|}K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}(r^{\prime},\theta^{\prime})\,\tfrac{e^{-i\ell\theta^{\prime}}}{\sqrt{2\pi}}\,.$

### v3

@@ 2.3. Spectral and scattering properties
Making reference to the quadratic form parametrization $H_{\mathrm{P}}^{(\beta)}$ of the Pauli Hamiltonian, using the bijection in Proposition 2.9 and noting that $L(0)=0$ (see (2.12)), we get
@@ 2.4. Symmetries and connection with the Dirac operator
$\|H_{\mathrm{D}}\bm{\psi}\|^{2}=\left\langle\bm{\psi}\left|H_{\mathrm{P}}\right|\bm{\psi}\right\rangle=Q^{(\mathrm{F})}_{\mathrm{P}}(\bm{\psi})\,,$
@@
$\mathscr{D}\big{(}\overline{H}_{\mathrm{D}}\big{)}=\mathscr{D}\big{[}Q^{(\mathrm{F})}_{\mathrm{P}}\big{]}\,.$ (2.61)
@@
For any $\alpha\in(0,1)$, the symmetric operator $H_{\mathrm{D}}$ admits a one-parameter family of self-adjoint extensions $H_{\mathrm{D}}^{(\gamma)}$, parametrized by $\gamma\in[0,2\pi)$ and given by
$\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}=\left\{\bm{\psi}=\bm{\phi}+\mu\left(\bm{\xi}_{+}+e^{i\gamma}\bm{\xi}_{-}\right)\,\Big{|}\,\bm{\phi}\in\mathscr{D}\big{[}Q^{(\mathrm{F})}_{\mathrm{P}}\big{]}\,,\;\mu\in\mathbb{C}\right\},$ (2.63)
$H_{\mathrm{D}}^{(\gamma)}\bm{\psi}=H_{\mathrm{D}}\bm{\phi}+i\mu(\bm{\xi}_{+}-i\bm{\xi}_{-})\,.$ (2.64)
Alternatively, as illustrated in [Pe06], the domain of the extensions $H_{\mathrm{D}}^{(\gamma)}$ can be described in terms of boundary conditions at the origin, as follows: given $\gamma\in[0,2\pi)$, one can define the linear functionals $c^{\uparrow,\downarrow}_{-\alpha}$, $c^{\uparrow,\downarrow}_{\alpha-1}$ on $\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$ as
$c^{s}_{-\alpha}(\bm{\psi})=\lim_{r\to 0^{+}}r^{\alpha}\,\langle\psi_{s}\rangle=\tfrac{1}{2\pi}\lim_{r\to 0^{+}}r^{\alpha}\int^{2\pi}_{0}\mathrm{d}\vartheta\,\psi_{s}(r,\vartheta)\,,$ (2.65)
$c^{s}_{\alpha-1}(\bm{\psi})=\lim_{r\to 0^{+}}r^{1-\alpha}\,\langle e^{i\vartheta}\psi_{s}\rangle=\tfrac{1}{2\pi}\lim_{r\to 0^{+}}r^{1-\alpha}\int^{2\pi}_{0}\mathrm{d}\vartheta\,e^{i\vartheta}\,\psi_{s}(r,\vartheta)\,,$ (2.66)
for $s\in\left\{\uparrow,\downarrow\right\}$. Observe that, for a given $\bm{\psi}\in\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$, we have a decomposition $\bm{\psi}=\bm{\phi}+\mu(\bm{\xi}_{+}+e^{i\gamma}\bm{\xi}_{-})$ as in Proposition 2.24, so that $\bm{\phi}$ vanishes at the origin, thus having zero boundary value. The asymptotics of Bessel functions $K_{\nu}$ at the origin [GR07] then easily give
@@
where $\Gamma(z)$ is the Euler gamma function. We can then equivalently describe the domain of $H_{\mathrm{D}}^{(\gamma)}$ as
$\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})=\left\{\bm{\psi}\in L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})\,\Big{|}\,H_{\mathrm{D}}\bm{\psi}\in L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})\,,\;c^{\uparrow}_{\alpha-1}(\bm{\psi})=i\cot(\gamma/2)\,\tfrac{2^{1-2\alpha}\Gamma(1-\alpha)}{\Gamma(\alpha)}\,c^{\downarrow}_{-\alpha}(\bm{\psi})\,,\right.\\ \left.c^{\uparrow}_{-\alpha}(\bm{\psi})=0\,,c^{\downarrow}_{\alpha-1}(\bm{\psi})=0\right\}.$ (2.69)
Remark 2.25 (Local singularities of functions in the Dirac domain).
For a generic $\gamma\in[0,2\pi)$, considering the characterization (2.63) of $\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}$, the asymptotic properties in (2.5), the expressions (2.62) of $\bm{\xi}_{\pm}$ and the expansions of the Bessel functions for small argument, for any $\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}$ one infers that, as $r\to 0^{+}$,
$\bm{\psi}(r,\theta)=\mu\begin{pmatrix}(1+e^{i\gamma})\,\frac{2^{-\alpha}\Gamma(1-\alpha)}{r^{1-\alpha}}\,e^{-i\vartheta}\vspace{0.15cm}\\ (1-e^{i\gamma})\,\frac{2^{-(1-\alpha)}\Gamma(\alpha)}{r^{\alpha}}\end{pmatrix}+o(1)=\tilde{\mu}\begin{pmatrix}i\cot(\gamma/2)\,\frac{2^{1-2\alpha}\Gamma(1-\alpha)}{\Gamma(\alpha)}\,\frac{1}{r^{1-\alpha}}\,e^{-i\vartheta}\vspace{0.15cm}\\ \frac{1}{r^{\alpha}}\end{pmatrix}+o(1)\,,$
for some $\tilde{\mu}\in\mathbb{C}$.
We have already observed that the symmetric operators $H_{\mathrm{P}}$ and $H_{\mathrm{D}}$ defined on $C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{0\};\mathbb{C}^{2})$, are such that $H_{\mathrm{D}}^{2}=H_{\mathrm{P}}$. Now we investigate whether the same property holds for the the respective self-adjoint extensions. More precisely, we aim at determining which extension $H_{\mathrm{P}}^{(\beta)}$ is the square of an extension $H_{\mathrm{D}}^{(\gamma)}$, if any. It turns out that the answer to this question is actually trivial, since the square of all self-adjoint realizations of the Dirac operator coincides with the Fridriechs extension of $H_{\mathrm{P}}$. As a byproduct of the proof, we are also able to classify all the extensions $H_{\mathrm{P}}^{(\beta)}$ such that $\mathscr{D}(H_{\mathrm{P}}^{(\beta)})\subseteq\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$ for some $\gamma\in[0,2\pi)$, i.e., both the first and second order operators are simultaneously well-posed, although in general the latter is not the square of the former.
@@
$\left(H_{\mathrm{D}}^{(\gamma)}\right)^{2}=H_{\mathrm{P}}^{(\mathrm{F})}\,.$ (2.70)
Furthermore, $\mathscr{D}(H_{\mathrm{P}}^{(\beta)})\subseteq\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$, for some $\gamma\in[0,2\pi)$, if and only if, for all $\lambda>0$ large enough,
@@
In passing, we point out the following result which has its own interest: unlike Pauli operators with an AB flux, any self-adjoint realization of the Dirac operator with an AB flux admits a zero-energy resonance.
Proposition 2.27 (Zero-energy resonances for $H_{\mathrm{D}}^{\gamma}$).
For any $\gamma\in[0,2\pi)$, the Dirac Hamiltonian $H_{\mathrm{D}}^{\gamma}$ has a zero-energy resonance $\bm{\psi}_{0}$ given by
$\bm{\psi}_{0}(r,\vartheta)=\begin{cases}\begin{pmatrix}i\cot(\gamma/2)\,\tfrac{2^{1-2\alpha}\Gamma(1-\alpha)}{\Gamma(\alpha)}\,e^{-i\vartheta}\,r^{-(1-\alpha)}\vspace{0.1cm}\\ r^{-\alpha}\end{pmatrix}\,&\mbox{for }\gamma\neq 0,\\ \begin{pmatrix}e^{-i\vartheta}\,r^{-(1-\alpha)}\\ 0\end{pmatrix},&\mbox{for }\gamma=0.\end{cases}$ (2.72)
Remark 2.28 (Resonances for Dirac and Pauli operators).
The above result is consistent with the identity $\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}^{2}=H^{(\mathrm{F})}_{\mathrm{P}}$, stated in Proposition 2.26, and with the fact that $H^{(\mathrm{F})}_{\mathrm{P}}$ has no zero-energy resonance (see Proposition 2.21). Indeed, for any given $\gamma\in[0,2\pi)$, even if the spinor $\bm{\psi}_{0}$ found in Proposition 2.27 does satisfy the distributional equation $H_{\mathrm{P}}\bm{\psi}_{0}=(H_{\mathrm{D}}^{\gamma})^{2}\bm{\psi}_{0}=0$, it is not a resonance for the Friedrichs extension $H^{(\mathrm{F})}_{\mathrm{P}}$. This is due to the singular behavior at the origin, making it impossible to satisfy the boundary conditions encoded in $\mathscr{D}(H_{\mathrm{P}}^{(\mathrm{F})})$ (see item $ii)$ in Proposition 2.1 and Definition 2.1).
@@ 3.1. Quadratic forms
$g^{(\ell)}_{\lambda}\!(r,\vartheta)=\left[\tfrac{\Gamma(|\ell+\alpha|)}{2^{1-|\ell+\alpha|}}\,{1\over r^{|\ell+\alpha|}}+\tfrac{\Gamma(\!-|\ell+\alpha|)}{2^{1+|\ell+\alpha|}}\,\lambda^{2|\ell+\alpha|}\,r^{|\ell+\alpha|}+\mathcal{O}\big{(}r^{2-|\ell+\alpha|}\big{)}\right]\tfrac{e^{i\ell\vartheta}}{\sqrt{2\pi}}\,,\quad\mbox{for\, $r\to 0^{+}$}.$ (3.2)
@@ 3.2. Krein’s theory
$\big{|}I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z}\,(r\land r^{\prime})\big{)}\big{|}\leqslant\left(\!\tfrac{\sqrt{|z|}}{\Im\sqrt{z}}\right)^{\!|\ell+\alpha|}I_{|\ell+\alpha|}\big{(}\Im\sqrt{z}\,(r\land r^{\prime})\big{)}\,;$ (3.5)
$\big{|}K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z}\,(r\lor r^{\prime})\big{)}\big{|}\leqslant K_{|\ell+\alpha|}\big{(}\Im\sqrt{z}\,(r\lor r^{\prime})\big{)}\,.$ (3.6)
@@
$\tau_{s}^{(\ell)}R_{\mathrm{S}}^{(\mathrm{F})}(z^{*})\,\bm{\psi}\\ =\lim_{r\to 0^{+}}\!\tfrac{2^{|\ell+\alpha|-1}\,\Gamma\big{(}|\ell+\alpha|\big{)}}{r^{|\ell+\alpha|}}\;\big{(}|\ell+\alpha|+r\,\partial_{r}\big{)}\!\int^{\infty}_{0}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\,I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,(r\land r^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,(r\lor r^{\prime})\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\\ =2^{|\ell+\alpha|-1}\,\Gamma\big{(}|\ell+\alpha|\big{)}\,i\sqrt{z^{*}}\lim_{r\to 0^{+}}\!\,r^{1-|\ell+\alpha|}\,\bigg{[}K_{|\ell+\alpha|-1}\big{(}\!-i\sqrt{z^{*}}\,r\big{)}\!\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\\ -I_{|\ell+\alpha|-1}\big{(}\!-i\sqrt{z^{*}}\,r\big{)}\!\int^{\infty}_{r}\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\,K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\bigg{]}\,.$
@@
$\left|\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}^{(\ell)}(r^{\prime})\right|\leqslant\big{\|}\psi_{s}^{(\ell)}\big{\|}_{L^{2}(\mathbb{R}^{+},\,r\,dr)}\left(\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,\big{|}I_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\big{|}^{2}\right)^{\!\!1/2}\\ \leqslant\left(\!\tfrac{\sqrt{|z|}}{\Im\sqrt{z^{*}}}\right)^{\!|\ell+\alpha|}\big{\|}\psi_{s}^{(\ell)}\big{\|}_{L^{2}(\mathbb{R}^{+},\,r\,\mathrm{d}r)}\left(\int^{r}_{0}\!\mathrm{d}r^{\prime}\,r^{\prime}\,I^{2}_{|\ell+\alpha|}\big{(}\Im\sqrt{z^{*}}\,(r\land r^{\prime})\big{)}\!\right)^{\!\!1/2}\leqslant C\,\|\bm{\psi}\|_{2}\;r^{1+|\ell+\alpha|}\xrightarrow[r\to 0^{+}]{}0\,,$
@@
$\tau_{s}^{(\ell)}R_{\mathrm{S}}^{(\mathrm{F})}(z^{*})\,\bm{\psi}=\int_{0}^{\infty}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\!\int_{0}^{2\pi}\!\!\!\!\mathrm{d}\vartheta^{\prime}\,\big{(}\!-i\sqrt{z^{*}}\big{)}^{|\ell+\alpha|}K_{|\ell+\alpha|}\big{(}\!-i\sqrt{z^{*}}\,r^{\prime}\big{)}\,\psi_{s}(r^{\prime},\vartheta^{\prime})\,\tfrac{e^{-i\ell\vartheta^{\prime}}}{\sqrt{2\pi}}\,.$